Which among the following statements is CORRECT? O Electric potential is a vector quantity 。If the strength of a uniform electric field doubles, then the associated electric potential difference must double f the charge on a capacitor is doubled, the energy stored in the capacitor must have halved OThe capacitance of a capacitor is determined by the size of the battery it is connected to. QUESTION 2 A charge is released from rest in a uniform electric field of magnitude 1000 V/m (Volt per mater, Find the change in displacement of 0.3 m (meter). electric potential of the charge after it undergoes a O -150 V -3000 V O-120 V -300 V
Answer #1

(1)

(a) Electric potential is scalar quantity.

(c) Energy stored in the capacitor is given by E = 1/2*Q*V

Thus, if charge is doubled, energy stored also gets doubled.

(d) Capacitance is the property of the capacitor and is independent of battery connected to it.

Since,

Potential difference, V = E*L where E is electric field and L is distance.

Thus, if Electric field is doubled corresponding Potential also gets doubled.

(2)

Given,

Magnitude of electric field, E = 1000 V/m

displacement, d = 0.3 m

Change in electric potential, V = -E*d = -1000*0.3

                                                  = -300 V